# AutoCheck: Automatically Identifying Variables for Checkpointing by Data Dependency Analysis

### v2

@@ I Introduction
To solve this problem, we propose AutoCheck, an analytical model and tool that can automatically identify critical variables (or variables necessary for checkpointing to correctly restart upon failures). Our solution is the first of its kind. AutoCheck is a complement to existing C/R libraries. AutoCheck frees programmers and domain scientists from prohibitive workloads, which require understanding algorithm semantics and manually tracking complicated data dependency across nested loops and function calls, via efficiently tracking dynamic data dependency and a collection of heuristics for identifying critical variables for checkpointing automatically.
AutoCheck’s analytical model takes the dynamic application execution trace as input and has three key modules. First, identifying the main computation loop’s input variables, for which we identify these variables defined before and used inside the main computation loop for further analysis. This allows us to confine the critical variables to analyze. Second, we track and extract the data dependency only to those variables by selectively iterating the dynamic instructions, and propose a composition of optimizations to construct a contracted data dependency graph (DDG). It allows us to quickly and progressively pinpoint useful information from complicated data dependencies. Third, for the first time, we summarize three data dependency patterns at individual variables for C/R, which are write-after-read, outcome, partially-overwritten.
@@
  2. 2.
    An analysis and formal definition of three key data dependency patterns at individual variables for C/R that we summarize from practices;
@@ III Complexity of Manual Inspection
(3) Convoluted data dependency: Taking as an example u(double***[5]) in BT, a 4D array dependent on as many as 17 other different variables across many distinct function invocations. AutoCheck addresses all the data dependency on u(double***[5]) and finds it involved in Write-After-Read, which requires checkpointing.
@@ IV-C Identification of critical variables
This module takes the contracted DDG as input and outputs critical variables. The contracted DDG is stored as a linked list, from which we pull out read and write dependencies on individual variables by converting the contracted DDG into an execution-time-ordered sequence of read and write dependencies, according to the order of each dynamic instruction in execution. We give an example of the extracted read and write dependencies in execution time order in Figure 5(e). With the obtained read and write dependencies, we can address the data dependency on individual variables to identify critical variables. We propose three dependency types to identify critical variables.
@@
Read-After-Partially-Overwritten (RAPO): When a variable is an array, a write to it can result in partial overwriting, where only some of its elements are updated before they are read by others. In the event of a failure during restart, the elements that were not involved in the overwriting cannot be recovered. To maintain the state of the array, it is necessary to checkpoint it. An example is ‘$a$’ (at Line 16) in the example code illustrated in Figure 5(d).
We select the variables for checkpointing according to the three dependency types mentioned above. AutoCheck performs element-level dependency analysis. Note that if a variable’s dependency pattern does not belong to any of the three types, there is no need to checkpoint.
Index. In addition, we also do checkpoint to the induction variables of the main computation loop (only the outermost loop of the main computation loop). And we use llvm-pass-loop API for identification.
This is a separate case. The reason is that in a restart, we can direct the execution to the iteration where it dropped out because of the failure. For example, as shown in Figure 4, Line 13 is the ‘for’ conditional statement of the main computation loop, in which ‘$it$’ is the induction variable to checkpoint.
@@ V-C Data dependency analysis
AMG (ECP) OMP+MPI Algebraic Multi-Grid linear system solver for unstructured mesh physics packages (input size $-problem$ 2 $-n$ $5\times 5\times 5$ 75000 6.8G 117.39 $diagonal$ (WAR), $cum\_num\_its$ (WAR), $cum\_nnz\_AP$ (WAR), $hypre\_\_global\_error$ (WAR), $final\_res\_norm$ (Outcome), $j$ (Index) 462-553 (amg.c)
HACC OMP+MPI The Hardware Accelerated Cosmology Code framework(input size $-N$ 1 $-t$ $2\times 2\times 2$ 32254 12.7G 201.13 particles (WAR),step (Index) 318-523 (driver_hires-local.cxx)
@@
AMG -problem 2 -n 40 40 40 647577.68 58.83
HACC -$N$ 8 -$t$ $16\times 16\times 16$ 837533.14 334.93
@@ VIII Related work
Checkpoint/Restart. There have been a variety of C/R libraries [13, 9, 10, 11, 12, 14, 30] developed with different focuses. Bautista-Gome et al. [13] developed Fault Tolerant Interface (FTI), which is a scalble, multi-level checkpointing library. Nicolae et al. [14] provided an asynchronous checkpointing solution. However, all those C/R libraries request programmers with little domain knowledge to identify variables to checkpoint manually. To solve this problem, we develop a tool, AutoCheck, that leverages data dependency analysis and a series of heuristics to automatically identify variables for checkpointing.
Data-dependency-based resilience analytic. Numerous existing research [31, 32, 33, 34] leverages data dependency analysis to do resilience analytics and to identify code vulnerabilities. Guo et al. [35] developed an analytical tool for measuring application vulnerability on individual variables by counting error masking events from data-dependency analysis. Menon et al. [36] leveraged Automatic Differentiation (AD) for data correlation analysis to identify resilient variables to transient faults. In contrast, our work aims to free programmers from manually tracking complicated data dependency across loops and functions to identify variables for checkpointing. Our work extends the capability of C/R libraries.

### v3

@@ I Introduction
To solve this problem, we propose AutoCheck, an analytical model and tool that can automatically identify critical variables (or the minimal set of variables necessary for checkpointing to correctly restart upon failures). Our solution is the first of its kind. AutoCheck is a complement to existing C/R libraries. AutoCheck frees programmers and domain scientists from prohibitive workloads, which require understanding algorithm semantics and manually tracking complicated data dependency across nested loops and function calls, via efficiently tracking dynamic data dependency and a collection of heuristics for identifying critical variables for checkpointing automatically.
AutoCheck’s analytical model takes the dynamic application execution trace as input and has two key modules. First, identifying the main computation loop’s input variables, for which we identify these variables defined before and used inside the main computation loop for further analysis. This allows us to confine the critical variables to analyze. Second, we track and extract the data dependency only to those variables by selectively iterating the dynamic instructions, and propose a composition of optimizations to construct a contracted data dependency graph (DDG). It allows us to quickly and progressively pinpoint useful information from complicated data dependencies.
@@
    A trace analysis tool that can automatically identify variables necessary for checkpointing to recover from system failures;
  2. 2.
@@ III Complexity of Manual Inspection
(3) Convoluted data dependency: Taking as an example u(double****) in BT, a 4D array dependent on as many as 17 other different variables across many distinct function invocations. AutoCheck addresses all the data dependency on u and finds it involved in Write-After-Read, which requires checkpointing.
@@ IV-C Identification of critical variables
This module takes the contracted DDG as input and outputs critical variables. The contracted DDG is stored as a linked list, from which we pull out read and write dependencies on individual variables by converting the contracted DDG into an execution-time-ordered sequence of read and write dependencies, according to the order of each dynamic instruction in execution. We give an example of the extracted read and write dependencies in execution time order in Figure 5(e). With the obtained read and write dependencies, we can address the data dependency on individual variables to identify critical variables. A recent study [20] groups critical variables to checkpoint into four types based on data dependency patterns.
@@
Read-After-Partially-Overwritten (RAPO): When a variable is an array, a write to it can result in partial overwriting, where only some of its elements are updated before they are read by others. In the event of a failure during restart, the elements that were not involved in the overwriting cannot be recovered. To maintain the state of the array, it is necessary to checkpoint it.
An example is ‘$a$’ (at Line 16) in the example code illustrated in Figure 5(d).
Index. We also do checkpoint to the induction variables of the main computation loop (only the outermost loop of the main computation loop). And we use llvm-pass-loop API for identification. This is a separate case. The reason is that in a restart, we can direct the execution to the iteration where it dropped out because of the failure. For example, as shown in Figure 4, Line 13 is the ‘for’ conditional statement of the main computation loop, in which ‘$it$’ is the induction variable to checkpoint.
@@ V-C Data dependency analysis
AMG (ECP) OMP+MPI Algebraic Multi-Grid linear system solver for unstructured mesh physics packages (input size $-problem$ 2 $-n$ $5\times 5\times 5$) 75000 6.8G 117.39 $diagonal$ (WAR), $cum\_num\_its$ (WAR), $cum\_nnz\_AP$ (WAR), $hypre\_\_global\_error$ (WAR), $final\_res\_norm$ (Outcome), $j$ (Index) 462-553 (amg.c)
HACC OMP+MPI The Hardware Accelerated Cosmology Code framework(input size $-N$ 1 $-t$ $1\times 1\times 1$) 32254 12.7G 201.13 particles (WAR),step (Index) 318-523 (driver_hires-local.cxx)
@@
AMG -problem 2 -n 40 40 40 647577.68 0.01
HACC -$N$ 1 -$t$ $1\times 1\times 1$ 217533.14 134.93
@@ VIII Related work
Checkpoint/Restart. There have been a variety of C/R libraries [13, 9, 10, 11, 12, 14, 31, 29, 32] developed with different focuses. Bautista-Gome et al. [13] developed Fault Tolerant Interface (FTI), which is a scalble, multi-level checkpointing library. Nicolae et al. [14] provided an asynchronous checkpointing solution. However, all those C/R libraries request programmers with little domain knowledge to identify variables to checkpoint manually. To solve this problem, we develop a tool, AutoCheck, that leverages data dependency analysis and a series of heuristics to automatically identify variables for checkpointing.
Data-dependency-based resilience analytic. Numerous existing research [33, 34, 35, 36, 37, 38, 39, 40, 41, 42] leverages data dependency analysis to do resilience analytics and to identify code vulnerabilities. Guo et al. [38] developed an analytical tool for measuring application vulnerability on individual variables by counting error masking events from data-dependency analysis. Menon et al. [43] leveraged Automatic Differentiation (AD) for data correlation analysis to identify resilient variables to transient faults. In contrast, our work aims to free programmers from manually tracking complicated data dependency across loops and functions to identify variables for checkpointing. Our work extends the capability of C/R libraries.